OpenStudy (anonymous):

A single strand of DNA helix consists of 100 nitrogenous base pairs. If 36 of the bases are adenine and 24 are cytosine, how many should be thymine on the complementary strand?

OpenStudy (anonymous):

Hi,well generally according to Chargaff low an SS DNA wich have 36 A the complementary strand should have 36 T , and according to complementary dictate number of purines equal pyrimidines. Chargaff low: (A+T) = (C+G) \[\rightarrow \frac{ A }{ T } = \frac{ C }{ G } =1\]. You can reading the chapter 19 principles of biochemistry Horton. Moran. Rawn or principles of genetics Tamarin. or molecular biology of the gene Mr.Watson. excuse me my English isn't well so maybe i have Grammatical errors.